# On sporadic symmetry breaking operators for principal series representations of the de Sitter and Lorentz groups

### v1

@@ 6 Localness Theorem
The first assertion is a direct consequence of (7.3). Let us prove the second one, in which the arguments are very similar to those of [20, Thm. 2.10]. First, note that
$G/G_{0}\simeq G^{\prime}/G_{0}^{\prime}\simeq({\mathbb{Z}}/2{\mathbb{Z}})^{2},$
$(G/G_{0})^{\wedge}\simeq(G^{\prime}/G_{0}^{\prime})^{\wedge}\simeq\{\chi_{a,b}:a,b=\pm\},$
where $\chi_{a,b}:G\rightarrow\{\pm\}$ are one-dimensional representations characterized by
$\chi_{a,b}(\text{{diag}}(-1,1,1,1,1))=a,$ $\chi_{a,b}(\text{{diag}}(1,1,1,1,-1))=b.$
By little abuse of notation, we use $\chi_{a,b}$ to denote representations of $G$ as well as of $G/G_{0}\simeq G^{\prime}/G_{0}^{\prime}$. Now, for
$p$ $=(N,\alpha,\lambda)\in{\mathbb{N}}\times{\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{C}},$
$q$ $=(m,\beta,\nu)\in{\mathbb{N}}_{+}\times{\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{C}},$
set
$\begin{gathered}I(p):=I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)_{\alpha},\quad J(q):=J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta},\\ S(p,q):=\text{{Hom}}_{G^{\prime}_{0}}(I(p),J(q)).\end{gathered}$
Let $G^{\prime}/G_{0}^{\prime}$ act on $S(p,q)$ by ${\mathbb{T}}\mapsto\varpi(g)({\mathbb{T}}):=J(q)(g)\circ{\mathbb{T}}\circ I(p)(g^{-1})$ for $gG_{0}^{\prime}\in G^{\prime}/G_{0}^{\prime}$. This is a well-defined action as $G_{0}^{\prime}$ is a normal subgroup of $G^{\prime}$. Moreover, since $S(p,q)$ is finite-dimensional (see [19, 24]), we have an irreducible decomposition
$S(p,q)\simeq\bigoplus_{\chi\in(G^{\prime}/G_{0}^{\prime})^{\wedge}}S(p,q)_{\chi},$ (7.5)
where $S(p,q)_{\chi}$ denotes the $\chi$-component of $S(p,q)$ given by
$S(p,q)_{\chi}=\{{\mathbb{T}}\in S(p,q):\varpi(g)({\mathbb{T}})=\chi(g){\mathbb{T}}\text{ for all }gG_{0}^{\prime}\in G^{\prime}/G_{0}^{\prime}\}.$
By the expression of $S(p,q)_{\chi}$ above, we obtain that
$S(p,q)_{\chi}\simeq\text{{Hom}}_{G^{\prime}}(I(p),J(q)\otimes\chi).$
Moreover, a straightforward computation shows that $J(q)\otimes\chi_{a,b}$ is given as follows (see, for instance [28, Lem. 2.14]):
$J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta}\otimes\chi_{++}=J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta},$
$J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta}\otimes\chi_{+-}=J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta+1},$
$J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta}\otimes\chi_{--}=J(\mathcal{H}^{m}({\mathbb{C}}^{2})\otimes\det,\nu)_{\beta},$
$J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\beta}\otimes\chi_{-+}=J(\mathcal{H}^{m}({\mathbb{C}}^{2})\otimes\det,\nu)_{\beta+1}.$
On the other hand, since $m>N$, we have
$[\mathcal{H}^{N}({\mathbb{C}}^{3}):\mathcal{H}^{m}({\mathbb{C}}^{2})]=\dim_{\mathbb{C}}\text{{Hom}}_{O(2)}(\mathcal{H}^{N}({\mathbb{C}}^{3})\big{\rvert}_{O(2)},\mathcal{H}^{m}({\mathbb{C}}^{2}))=0,$
as the branching law for $\mathcal{H}^{N}({\mathbb{C}}^{3})$ with respect to $O(3)\downarrow O(2)$ is given by (cf. [28, Fact 2.12]):
$\mathcal{H}^{N}({\mathbb{C}}^{3})\big{\rvert}_{O(2)}=\mathcal{H}^{N}({\mathbb{C}}^{3})\big{\rvert}_{O(2)}\simeq\bigoplus_{k=\varepsilon}^{N}\mathcal{H}^{k}({\mathbb{C}}^{2}),\qquad\varepsilon=N\mod 2.$
Therefore, by [28, Thms. 3.6 and 5.21], we have
$\text{{Hom}}_{G^{\prime}}(I(\mathcal{H}^{N}({\mathbb{C}}^{3}),\lambda)_{\gamma},J(\mathcal{H}^{m}({\mathbb{C}}^{2}),\nu)_{\delta})=\{0\},$
for any $\gamma,\delta\in{\mathbb{Z}}/2{\mathbb{Z}}$ with $\gamma-\delta\not\equiv\nu-\lambda\mod 2$. Hence, the right-hand side of (7.5) has only two summands, the ones corresponding to $\chi_{++}$ and $\chi_{--}$. The decomposition (7.4) follows now from (7.5). ∎
Now, we are ready to prove Theorem 7.3.
Proof of Theorem 7.3.
Take $(N_{1},m_{1})\in{\mathbb{N}}\times{\mathbb{Z}}$ with $|m_{1}|>N_{1}$ and denote by $\mathcal{P}_{0}$ the set of $(\lambda,\nu)\in{\mathbb{C}}^{2}$ such that (1.1) is non-zero. We take $(\lambda_{1},\nu_{1})\in\mathcal{P}_{0}$ and show that any SBO in (1.1) for $(N_{1},m_{1},\lambda_{1},\nu_{1})$ is sporadic with respect to the pair $(G_{0},G_{0}^{\prime})$. By definition, this is to find an open neighbourhood $U_{1}$ of $(\lambda_{1},\nu_{1})$ such that $U_{1}\cap\mathcal{P}_{0}=\{(\lambda_{1},\nu_{1})\}$.
Set $V_{1}:=\mathcal{H}^{N_{1}}({\mathbb{C}}^{3})\in\widehat{O(3)}$, $W_{1}:=\mathcal{H}^{|m_{1}|}({\mathbb{C}}^{2})\in\widehat{O(2)}$ and take $\alpha_{1},\beta_{1}\in{\mathbb{Z}}/2{\mathbb{Z}}$ with $\beta_{1}-\alpha_{1}\equiv\nu_{1}-\lambda_{1}\mod 2$. Denote by $\mathcal{P}$ the set of $(\lambda,\nu)$ such that (7.1) is non-zero. We start by showing the following claim: Any SBO in (7.1) for the data $(V_{1},W_{1},\alpha_{1},\beta_{1},\lambda_{1},\nu_{1})$ is sporadic with respect to $(G,G^{\prime})$. In fact, first notice that by Theorems 1.3 and 1.6, the space (1.1) for the tuple $(N_{1},m_{1},\lambda_{1},\nu_{1})$ is spanned by $\mathbb{D}_{\lambda_{1},\nu_{1}}^{N_{1},m_{1}}$. Then, by Proposition 7.5(1), we have:
$\text{{Hom}}_{G_{0}^{\prime}}(I_{0}(V_{1},\lambda_{1}),J_{0}(W_{1},\nu_{1}))\simeq{\mathbb{C}}\mathbb{D}_{\lambda_{1},\nu_{1}}^{N_{1},m_{1}}\oplus{\mathbb{C}}\mathbb{D}_{\lambda_{1},\nu_{1}}^{N_{1},-m_{1}}.$
Therefore, and since $W_{1}\otimes\det\simeq W_{1}$ as $O(2)$-modules (cf. [28, Lem. 2.9]), by Proposition 7.5(2) we obtain
$\text{{Hom}}_{G^{\prime}}(I(V_{1},\lambda_{1})_{\alpha_{1}},J(W_{1},\nu_{1})_{\beta_{1}})\oplus\text{{Hom}}_{G^{\prime}}(I(V_{1},\lambda_{1})_{\alpha_{1}},J(W_{1},\nu_{1})_{\beta_{1}})\simeq{\mathbb{C}}\mathbb{D}_{\lambda_{1},\nu_{1}}^{N_{1},m_{1}}\oplus{\mathbb{C}}\mathbb{D}_{\lambda_{1},\nu_{1}}^{N_{1},-m_{1}}.$
In particular, (7.1) for $(V_{1},W_{1},\alpha_{1},\beta_{1},\lambda_{1},\nu_{1})$ is non-zero. This shows $\mathcal{P}_{0}\subset\mathcal{P}$. Now, as in the proof of Proposition 7.5, we deduce that
$[V_{1}:W_{1}]=[\mathcal{H}^{N_{1}}({\mathbb{C}}^{3}):\mathcal{H}^{|m_{1}|}({\mathbb{C}}^{2})]=\dim_{\mathbb{C}}\text{{Hom}}_{O(2)}(\mathcal{H}^{N_{1}}({\mathbb{C}}^{3})\big{\rvert}_{O(2)},\mathcal{H}^{|m_{1}|}({\mathbb{C}}^{2}))=0,$
as $|m_{1}|>N_{1}$. Thus, by Proposition 7.4, any SBO in (7.1) for $(V_{1},W_{1},\alpha_{1},\beta_{1},\lambda_{1},\nu_{1})$ is sporadic with respect to the pair $(G,G^{\prime})$. By definition, this means that there is an open neighbourhood $U_{1}$ of $(\lambda_{1},\nu_{1})\in{\mathbb{C}}^{2}$ such that $U_{1}\cap\mathcal{P}=\{(\lambda_{1},\nu_{1})\}$. This implies $U_{1}\cap\mathcal{P}_{0}\subset U_{1}\cap\mathcal{P}=\{(\lambda_{1},\nu_{1})\}$. In other words, any SBO in (1.1) for $(N_{1},m_{1},\lambda_{1},\nu_{1})$ is sporadic with respect to $(G_{0},G_{0}^{\prime})$. Now the theorem follows. ∎
@@ 8.1 Hypergeometric Series
Given $p,q\in{\mathbb{N}}$ and $a_{1},\ldots,a_{p},b_{1},\ldots,b_{q}\in{\mathbb{C}}$, we define the following series
@@
Here, $b_{j}$ are not zero nor negative integers, as that would make the denominator zero. The series above is known as the hypergeometric series, and its radius of convergence is known. For instance, if any $a_{j}$ is a non-positive integer, the series has a finite number of terms, and therefore it converges. In the general case when $a_{j}$ are not non-positive integers, by using the ratio test one can show that it converges absolutely for any $z\in{\mathbb{C}}$ if $p<q+1$, for $|z|<1$ if $p=q+1$, and that it diverges if $p>q+1$ (cf. [1, Thm. 2.1.1]). When this series has a non-zero radius of convergence, it defines an analytic function which is often called the generalized hypergeometric function. When $p=q+1=2$, this function is known as the ordinary (or Gaussian) hypergeometric function.
One of the most interesting and studied cases for $p=q+1$ is when $|z|=1$, since the ratio test does not give a general criterion of convergence. It can be proved, however, that the hypergeometric series ${}_{q+1}F_{q}$ converges absolutely for $|z|=1$ if $\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})>0$, conditionally for $z=e^{it}\neq 1$ if $-1<\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})\leq 0$ and diverges if $\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})\leq-1$. Also, several formulas are known for $z=1$, as is the Gauss formula for ${}_{2}F_{1}$:
Lemma 8.1 ([13, Eq. (48)]).
@@
Similar identities hold for greater $(p,q)$. One example is the Pfaff–Saalschütz identity for the function ${}_{3}F_{2}$:
@@
Some other examples are Dixon’s identity (see [1, Eq. (2.2.11)]), Thomae’s identity (see [1, Cor. 3.3.6]), or Sheppard’s identity (see [1, Cor. 3.3.4]) for ${}_{3}F_{2}$, or Dougall’s identities for ${}_{5}F_{4}$ and ${}_{7}F_{6}$ (see [1, Eqs. (2.2.9), (2.2.10)]).
@@ 8.2 Renormalized Gegenbauer polynomials
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the Gegenbauer polynomial (also known as ultraspherical polynomial), named after the Austrian mathematician Leopold H. Gegenbauer ([14]) is defined as follows (cf. [1, Sec. 6.4], [8, Sec. 3.15.1]):
$C_{\ell}^{\mu}(z)$ $:=\sum_{k=0}^{[\frac{\ell}{2}]}(-1)^{k}\frac{\Gamma(\ell+\mu-k)}{\Gamma(\mu)k!(\ell-2k)!}(2z)^{\ell-2k}$
@@
In ${\mathbb{C}}[z]\setminus\text{{Pol}}_{\ell}[z]_{\text{{{even}}}}$, the Gegenbauer differential equation may have other solutions depending on the parameters $\mu,\ell$. For a concrete statement of this fact see [26, Thm. 11.4].
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the imaginary Gegenbauer differential operator $S_{\ell}^{\mu}$ is defined as follows (cf. [20, (4.7)]):
@@
Next, we give some condition on the parameter $\mu\in{\mathbb{C}}$ for which the “higher terms” of $\widetilde{C}_{\ell}^{\mu}$ vanish. At the end, we recall a “degree decay” property that happens when special parameters are considered.
@@
Let $\mu\in{\mathbb{C}}$ and $\ell\in{\mathbb{N}}$. Then, the term $z^{\ell-2k}\enspace(0\leq k<\left[\frac{\ell}{2}\right])$ in $\widetilde{C}_{\ell}^{\mu}(z)$ vanishes if and only if $\mu\in M_{\ell}^{k}$. In particular, if $\mu\notin{\mathbb{Z}}$ then $\deg(\widetilde{C}_{\ell}^{\mu}(z))=\ell$.
@@ 8.3 Algebraic properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators
which coincides with the three-term relation [20, Eq. (14.15)]. Now suppose that (8.15) is true for some $d\in{\mathbb{N}}$. Then, for $d+1$, we have
@@
$\quad+\sum_{s=0}^{d}\begin{pmatrix}d\\ s\end{pmatrix}\frac{\Gamma\left(\mu+\ell+s-d\right)}{\Gamma\left(\mu+\ell-d\right)}\widetilde{C}_{\ell+2(s-d-1)}^{\mu-s}(it),$
@@
We remark that identities (8.17) and (8.18) above are a simple generalization of [20, Eq. (14.9)] and [20, Eq. (14.12)] respectively.

### v2

@@ 6 Localness Theorem
The proof is straightforward from Remark 7.2(1), that shows in particular that $\mathcal{P}$ is discrete. ∎
@@ 8.1 Hypergeometric Series
Given $p,q\in{\mathbb{N}}$ and $a_{1},\ldots,a_{p}\in{\mathbb{C}}$, $b_{1},\ldots,b_{q}\in{\mathbb{C}}\setminus{\mathbb{Z}}_{\leq 0}$, the hypergeometric series is defined as
@@
The radius of convergence of (8.1) is known. For instance, if some $a_{j}$ is a non-positive integer, the series has a finite number of terms. If $a_{j}$ are not non-positive integers, it converges absolutely for any $z\in{\mathbb{C}}$ if $p<q+1$, for $|z|<1$ if $p=q+1$, and it diverges if $p>q+1$ (cf. [1, Thm. 2.1.1]). When this series has a non-zero radius of convergence, it defines an analytic function which is often called the generalized hypergeometric function. When $p=q+1=2$, this function is known as the ordinary (or Gaussian) hypergeometric function.
When $p=q+1$ and $|z|=1$ there is not a general criterion of convergence, but it is known that $_{q+1}F_{q}$ converges absolutely for $|z|=1$ if $\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})>0$, conditionally for $z=e^{it}\neq 1$ if $1<\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})\leq 0$ and diverges if $\text{{Re}}(\sum_{j=1}^{q}b_{j}-\sum_{j=1}^{q+1}a_{j})\leq-1$. When $z=1$, several formulas giving the precise value of the series are known, as is the Gauss formula for $_{2}F_{1}$:
Lemma 8.1 ([15, Eq. (48)]).
@@
For $(p,q)=(2,3)$ we have the Pfaff–Saalschütz identity:
@@ 8.2 Renormalized Gegenbauer polynomials
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the Gegenbauer polynomial (also known as ultraspherical polynomial), named after the Austrian mathematician Leopold H. Gegenbauer ([16]) is defined as follows (cf. [1, Sec. 6.4], [10, Sec. 3.15.1]):
$C_{\ell}^{\mu}(z)$ $=\sum_{k=0}^{[\frac{\ell}{2}]}(-1)^{k}\frac{\Gamma(\ell+\mu-k)}{\Gamma(\mu)k!(\ell-2k)!}(2z)^{\ell-2k}$
@@
In ${\mathbb{C}}[z]\setminus\text{{Pol}}_{\ell}[z]_{\text{{{even}}}}$, the Gegenbauer differential equation may have other solutions depending on the parameters $\mu,\ell$. For a concrete statement of this fact see [29, Thm. 11.4].
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the imaginary Gegenbauer differential operator $S_{\ell}^{\mu}$ is defined as follows (cf. [23, (4.7)]):
@@
Next, we give a condition on the parameter $\mu\in{\mathbb{C}}$ for which the “higher terms” of $\widetilde{C}_{\ell}^{\mu}$ vanish. At the end, we recall a “degree decay” property that happens when special parameters are considered.
@@
Let $\mu\in{\mathbb{C}}$ and $\ell\in{\mathbb{N}}$. Then, the term $z^{\ell-2k}\kern 4.62497pt(0\leq k<\left[\frac{\ell}{2}\right])$ in $\widetilde{C}_{\ell}^{\mu}(z)$ vanishes if and only if $\mu\in M_{\ell}^{k}$. In particular, if $\mu\notin{\mathbb{Z}}$ then $\deg(\widetilde{C}_{\ell}^{\mu}(z))=\ell$.
@@ 8.3 Algebraic properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators
which coincides with the three-term relation [23, Eq. (14.15)]. Now suppose that (8.15) is true for some $d\in{\mathbb{N}}$. Then, for $d+1$, we have
@@
$\qquad+\sum_{s=0}^{d}\begin{pmatrix}d\\ s\end{pmatrix}\frac{\Gamma\left(\mu+\ell+s-d\right)}{\Gamma\left(\mu+\ell-d\right)}\widetilde{C}_{\ell+2(s-d-1)}^{\mu-s}(it),$
@@
We remark that identities (8.17) and (8.18) above are a simple generalization of [23, Eq. (14.9)] and [23, Eq. (14.12)] respectively.
@@ Acknowledgements
The author is grateful to Toshiyuki Kobayashi and Toshihisa Kubo for their insightful comments on this work. He also thanks Temma Aoyama for fruitful discussions regarding the use of hypergeometric series in the proof of Proposition 4.13. Finally, he extends his appreciation to the anonymous referee for a careful review and helpful comments on the manuscript.